What values for x must be excluded in the following fraction?
x+2/x^2-9x+14
Thank you.

Answer by sdmmadam@yahoo.com(530)

(x+2)/(x^2-9x+14)
=(x+2)/[(x-7)(x-2)] (by factorisation of the dr)
Note: If the problem is
(x-2)/(x^2-9x+14)
=(x-2)/[(x-7)(x-2)]
=1/(x-7) (canceling (x-2) )
Note: (x^2-9x+14) is a quadratic expression
and the middle term is splittable into two parts
(-7x) and (-2x) so that their sum is the middle term
and their product is the product of the square term and the constant term.
Here (-9x) = (-7x)+(-2x) and (-7x)X(-2x) = 14x^2 = (x^2)X(14)
(x^2-9x+14)
=x^2-7x-2x+14
=(x^2-7x)-2x+14
=x(x-7)-2(x-7)
=xp-2p where p = (x-7)
=p(x-2)
=(x-7)(x-2)